stacey would like to have 1 million available to her at retirement. Her investments have an average annual return of 12%.If she contributes $210 per month, will she reach her goal when she retires at 33 years
Answer #1

use financial calculator as below

FV=1000000 as the future value at retirement

PV=0 as present value

PMT=-120 as monthly contribution

I/Y=12/12=1 as the monthly interest rate

Click CPT

Click N=445.69 months or 37.14 years will take to reach 1 million amount from now, so she will not her goal when she retires at 33 years

the above is answer..